# Analysis of mean field games via Fokker-Planck-Kolmogorov equations:

### v1

@@ 3. Auxiliary results
$\mathbb{B}_{n}^{\varepsilon}(x,y,t)=\left(\begin{array}[]{l}b(x,t,\sigma^{n})+Q(x,t,\sigma^{n})u_{\varepsilon}(y,t)\\ b_{\varepsilon}(y,t,\sigma)+(Qu)_{\varepsilon}(y,t,\sigma)-y\frac{\varepsilon\phi(y)}{\mu_{t}^{\varepsilon}(y)}\\ \end{array}\right).$
@@
The last estimate and the convergence of measures $\mu_{t}^{\varepsilon}(y)\,dy$ to the measure $\mu_{t}$ as $\varepsilon\to 0$ imply that for every $t$ the sequence $\pi_{t}^{n,\varepsilon_{m}}$ is tight. It follow that for every $t$ one can extract a sequence $\varepsilon_{m}\to 0$ such that the measures $\pi_{t}^{n,\varepsilon_{m}}$ converge weakly to a probability measure. Let $\psi\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ and the support of $\psi$ be in $B\times B$, where $B$ is an open ball of radius $r$ centered at zero. Denote by $B^{\prime}$ the open ball of the radius $r+1$ centered at zero. Set
@@
where the function $|u|_{\varepsilon}$ is defined by the same formula as the function $u_{\varepsilon}$ where $u$ is replaced by $|u|$. Note that for a nonnegative number $C(B^{\prime},\psi)$ depending on $C(B^{\prime})$ and $\psi$ the estimate
@@
where $s,t\in[0,T^{\prime}]$ and $\omega_{\psi}(v)=C(B^{\prime},\psi)v+C(B^{\prime},\psi)vh^{-1}(R/2v)$. Repeating the arguments from Lemma 3.2, we extract a sequence $\varepsilon_{m}\to 0$ such that for every $t$ the measures $\pi^{n,\varepsilon_{m}}_{t}$ converge weakly to a probability measure $\pi^{n}_{t}$. Moreover, the mapping $t\mapsto\pi^{n}_{t}$ is continuous with respect to the weak topology. The passage to the limit in the integral identity defining the solution $\pi^{n,\varepsilon_{m}}_{t}$ is based on the following two observations. First, we note that the functions $a^{ij}$, $b^{i}$, $q^{im}$ are locally bounded and continuous in $x$. Second, let us consider a function $v$ such that $v$ is integrable on $\mathbb{R}^{d}\times[0,T]$ with respect to the measure $\mu_{t}(dy)\,dt$. Let $\widetilde{v}$ be a smooth function with a compact support on $\mathbb{R}^{d}\times[0,T]$. Set
$v_{\varepsilon}(y,t)=\frac{1}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}v(z,s)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds$
@@
$\widetilde{v}_{\varepsilon}(y,t)=\frac{1}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}\widetilde{v}(z,s)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds.$
@@
By choosing $\widetilde{v}$, one can make the right-hand side arbitrary small. Furthermore, the functions $\widetilde{v}_{\varepsilon}$ converge uniformly to the function $\widetilde{v}$ on $\mathbb{R}^{d}\times[0,T^{\prime}]$ as $\varepsilon\to 0$. Thus to prove the passage to the limit as $\varepsilon\to 0$ we replace the functions $a^{ij}$, $b^{i}$, $q^{im}u$ and $u$ in the expressions $a^{ij}_{\varepsilon}$, $b^{i}_{\varepsilon}$, $(Qu)_{\varepsilon}^{i}$, $u_{\varepsilon}$ by functions with compact supports and then we use the uniform convergence of these new expressions as $\varepsilon\to 0$. Similarly the passage to the limit is proved in [15].
@@
Let denote by $P_{R}^{\omega}$ the set of all measures $\Pi\in P_{R}$ such that the projection of $\Pi$ on $(x,t)$ belongs to the set $\mathcal{M}_{R,M}^{\omega}(V)$. Since $P_{R}$ and $\mathcal{M}_{R,M}^{\omega}(V)$ are compact sets in the weak topology, the set $P_{R}^{\omega}$ is compact in the weak topology.
Let us consider the mapping $\mathcal{F}$ from $\mathcal{M}_{R,M}^{\omega}(V)\times P_{R}^{\omega}$ to $\mathbb{R}$ given by the formula
$\mathcal{F}(\sigma,\Pi)=\int_{U\times\mathbb{R}^{d}\times[0,T]}f(u,x,t,\sigma)\Pi(dudxdt)+\int_{\mathbb{R}^{d}}g(x,\sigma)\,\mu_{T}(dx),$
@@
holds for every $t\in[0,T]$. Applying Fatou’s lemma, we derive the estimate
@@
$\int_{U\times[0,T]\times\mathbb{R}^{d}}(1-\zeta_{N}(u))|f(u,x,t,\sigma^{n})|\,\Pi^{n}(dudtdx)\leq\\ \int_{|u|>N}\Bigl{(}C_{h}h(|u|)+m_{\varepsilon}\Bigr{)}\Lambda(dudt)+\varepsilon TRe^{MT}.$
@@
$\int_{|u|>N}\Bigl{(}C_{h}h(|u|)+m_{\varepsilon}\Bigr{)}\Lambda(dudt)\leq\varepsilon.$
@@
$\int_{U\times\mathbb{R}^{d}\times[0,T]}(1-\eta_{k}(x))\zeta_{N}(u)|f(u,x,t,\sigma^{n})|\,\Pi^{n}(dudxdt)\leq\\ \leq\widetilde{m}_{\varepsilon}\Pi^{n}\bigl{(}\{(u,x,t)\colon|x|>k\}\bigr{)}+\varepsilon TRe^{MT}.$
@@
$\widetilde{m}_{\varepsilon}(\Pi^{n}+\Pi)\bigl{(}\{(u,x,t)\colon|x|>k\}\bigr{)}<\varepsilon$
@@
$\Bigl{|}\int_{U\times\mathbb{R}^{d}\times[0,T]}f(u,x,t,\sigma^{n})\,\Pi^{n}(dudxdt)-\int_{U\times\mathbb{R}^{d}\times[0,T]}\eta_{k}(x)\zeta_{N}(u)f(u,x,t,\sigma^{n})\,\Pi^{n}(dudxdt)\Bigr{|}\leq\\ 2\bigl{(}\varepsilon+\varepsilon TRe^{MT}\bigr{)}.$
@@
Let $\sigma\in\mathcal{M}_{R,M}(V)$ and $u_{0}\in U$. According to the condition (H2.1), we have the estimate
$L_{\sigma,u_{0}}V(x)\leq C_{L}V(x)+C_{L}Re^{Mt}+\beta_{V,W}(Re^{MT})Re^{MT}+h(|u_{0}|).$
By Remark 3.4 there exists a probability solution $\mu_{t}\,dt$ of the Cauchy problem
@@
Furthermore, for every $t\in[0,T]$ one has the estimate
$\int_{\mathbb{R}^{d}}V\,d\mu_{t}\leq e^{C_{L}T}\Bigl{(}\|V\|_{L^{1}(\nu)}+T\beta_{V,W}(Re^{MT})Re^{MT}+Th(|u_{0}|)\Bigr{)}+\frac{1}{4}Re^{Mt}.$
By Remark 3.2 we have $\lim_{R\to\infty}\beta_{V,W}(R)=0$. It follows that the estimate
$e^{C_{L}T}\Bigl{(}\|V\|_{L^{1}(\nu)}+T\beta_{V,W}(Re^{MT})Re^{MT}+Th(|u_{0}|)\Bigr{)}\leq\frac{3R}{4}$
holds for a sufficiently large $R>0$. Hence for every $t\in[0,T]$ we have
$\int_{\mathbb{R}^{d}}V\,d\mu_{t}\leq Re^{Mt}.$
Set
@@
Applying the conditions (H3.1) and (H3.2), we obtain the estimate
@@
Thus it suffices to minimize the functional $(u,\mu)\mapsto F_{\sigma}(u,\mu)$ only on the set of pairs $(u,\mu)$ such that
@@
Note that the right–hand side has the form $\alpha(Re^{MT})Re^{MT}$, where $\lim_{R\to\infty}\alpha(R)=0$. By Lemma 3.1 there exists a number $R_{0}>0$ such that for all $R>R_{0}$ the inequalities
$\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}V\,d\mu_{t}\leq Re^{Mt},\quad\int_{0}^{T}\int_{\mathbb{R}^{d}}h(|u(x,t)|)\,d\mu_{t}\,dt\leq\gamma R$
are fulfilled for every Borel mapping $u\colon\mathbb{R}^{d}\times[0,T]\to U$ and for every measure $\mu\in\mathcal{M}(V)$ satisfying the following conditions: 1) the measure $\mu=\mu_{t}\,dt$ is a solution to the Cauchy problem $\partial_{t}\mu=L^{*}_{\sigma,u(x,t)}\mu$, $\mu_{0}=\nu$, 2) the estimate $F_{\sigma}(u,\mu)\leq\alpha(Re^{MT})Re^{MT}$ holds. Taking into account Remark 3.7, we conclude that it suffices to minimize the functional $(u,\mu)\mapsto F_{\sigma}(u,\mu)$ only on the set of pairs $(u,\mu)$ such that
@@
on $U\times\mathbb{R}^{d}\times[0,T]$. If $\mu$ is a solution to the Cauchy problem $\partial_{t}\mu_{t}=L_{\sigma,u(x,t)}^{*}\mu_{t}$, $\mu_{0}=\nu$, then $\Pi$ belongs to the set $S_{R}(\sigma)$.
@@
$\mathcal{F}(\sigma,\Pi)=\int_{U\times\mathbb{R}^{d}\times[0,T]}f(u,x,t,\sigma)\Pi(dudxdt)+\int_{\mathbb{R}^{d}}g(x,\sigma)\,\mu_{T}(dx).$
@@
the measure $\widetilde{\Pi}$ belongs to the set $S_{R}(\sigma)$. Thus we can minimize the functional $\mathcal{F}_{\sigma}(\Pi)$ on $S_{R}(\sigma)$ instead of $F_{\sigma}(u,\mu)$. Let us take a number $R>0$ such that for $S_{R}(\sigma)$ the statements of Lemma 3.3 and Lemma 3.4 hold.
Denote by $M_{R}(\sigma)$ the set of all minimizers of the functional $\mathcal{F}_{\sigma}(\Pi)$ on $S_{R}(\sigma)$. By Lemma 3.5 the mapping $\Pi\mapsto\mathcal{F}_{\sigma}(\Pi)$ is lower semi-continuous. By Lemma 3.2 the set $S_{R}(\sigma)$ is compact. Thus $M_{R}(\sigma)$ is compact and contains at least one element. Moreover, the set $M_{R}(\sigma)$ is convex since the mapping $\Pi\mapsto\mathcal{F}_{\sigma}(\Pi)$ is linear.
@@
To prove Theorem 2.1 it suffices to find a measure $\mu\in\mathcal{M}_{R,M}^{\omega}(V)$ such that there exists a minimizer $\Pi$ of the functional $\mathcal{F}_{\mu}(\Pi)$ on the set $S_{R}(\mu)$ and the projection of $\Pi$ on $(x,t)$ is equal to the measure $\mu$. Thus the measure $\mu$ is a fixed point of the mapping $\sigma\mapsto\Phi(\sigma)$, where $\Phi(\sigma)$ is the set of all projections of measures $\Pi\in M_{R}(\sigma)$ on $(x,t)$. Let $e_{x,t}(u,x,t)=(x,t)$. Denote by $\Pi\circ e_{x,t}^{-1}$ the projection of $\Pi$ on $(x,t)$. Note that the mapping $\Pi\mapsto\Pi\circ e_{x,t}^{-1}$ is linear and continuous with respect to the weak topology. It follows that the set $\Phi(\sigma)$ is nonempty, convex and compact in the weak topology. Recall that $\mathcal{M}_{R,M}^{\omega}(V)$ is a convex set.
The existence of a fixed point of $\Phi$ is based on the Kakutani–Ky Fan theorem (see, for instance [18]): if a multivalued mapping $\Psi$ from a convex compact set $\mathcal{C}$ in a locally convex space to the set of non-empty convex compact subsets of $\mathcal{C}$ has a closed graph, then there exists a point $p$ such that $p\in\Psi(p)$.
Thus we need to verify that our mapping $\Phi$ has a closed graph. Assume that $\mu^{n}\in\Phi(\sigma^{n})$, $\sigma^{n}\to\sigma$ and $\mu^{n}\to\mu$. Let $\mu^{n}=\Pi^{n}\circ e_{x,t}^{-1}$, where $\Pi^{n}\in M_{R}(\sigma^{n})$. Since $\Pi^{n}\in P_{R}^{\omega}$ and $P_{R}^{\omega}$ is a compact set, there exists a subsequence $\Pi^{n_{k}}$ such that measures $\Pi^{n_{k}}$ converge weakly to a measure $\Pi\in P_{R}^{\omega}$. Note that $\mu=\Pi\circ e_{x,t}^{-1}$. In what follows we assume that the original sequence $\Pi^{n}$ converges weakly to $\Pi$. We need to prove that $\Pi\in M_{R}(\sigma)$.
@@
Arguing as above, we get $\widetilde{\Gamma}\in S_{R}(\sigma)$ and $\mathcal{F}_{\sigma}(\widetilde{\Gamma})\leq\mathcal{F}_{\sigma}(\Gamma)$.
@@
Let us consider the function $u$ and the measure $\mu$ constructed in Theorem 2.1. The measure $\mu$ is a probability solution to the Cauchy problem $\partial_{t}\mu_{t}=L_{\mu,u(x,t)}^{*}\mu_{t}$, $\mu_{0}=\nu$. Moreover, we have $V\in L^{1}(\mu)$. According to the condition (H2.3), the coefficients of $L_{\mu,u(x,t)}$ are integrable on $\mathbb{R}^{d}\times[0,T]$ with respect to the measure $\mu_{t}\,dt$. Applying the superposition principle (see, for instance [42], [15]), we obtain a probability measure $P$ on the space $C([0,T],\mathbb{R}^{d})$ such that $\mu_{t}=P\circ e_{t}^{-1}$, where $e_{t}(\omega)=\omega(t)$, and for every function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ the process
@@
is a martingale with respect to the measure $P$ and the natural filtration $\mathcal{F}_{t}=\sigma(\omega(s),s\leq t)$. By Proposition 2.1 from [34, Chapter 4] there exists a filtered probability space $(\Omega,\mathcal{F}_{t},\mathcal{P})$ supporting a $\mathcal{F}_{t}$–Brownian motion $W$ and a $\mathcal{F}_{t}$–adapted process $X$ such that
@@
and $\mathcal{P}\circ X_{t}^{-1}=\mu_{t}$ for all $t\in[0,T]$.

### v2

@@ 3. Auxiliary results
$\mathbb{B}_{n}^{\varepsilon}(x,y,t)=\left(\begin{array}[]{l}b(x,t,\sigma^{n})+Q(x,t,\sigma^{n})u_{\varepsilon}(y,t)\\ b_{\varepsilon}(y,t,\sigma)+(Qu)_{\varepsilon}(y,t,\sigma)-y\displaystyle\frac{\varepsilon\phi(y)}{\mu_{t}^{\varepsilon}(y)}\\ \end{array}\right).$
@@
The last estimate and the convergence of measures $\mu_{t}^{\varepsilon}(y)\,dy$ to the measure $\mu_{t}$ as $\varepsilon\to 0$ imply that for every $t$ the sequence $\pi_{t}^{n,\varepsilon_{m}}$ is tight. It follow that for every $t$ one can find a subsequence $\varepsilon_{m}\to 0$ such that the measures $\pi_{t}^{n,\varepsilon_{m}}$ converge weakly to a probability measure. Let $\psi\in C_{0}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ and the support of $\psi$ be in $B\times B$, where $B$ is an open ball of radius $r$ centered at zero. Denote by $B^{\prime}$ the open ball of the radius $r+1$ centered at zero. Set
@@
where the function $|u|_{\varepsilon}$ is defined by the same formula as the function $u_{\varepsilon}$ replacing $u$ by $|u|$. Note that for a nonnegative number $C(B^{\prime},\psi)$ depending on $C(B^{\prime})$ and $\psi$ the estimate
@@
where $s,t\in[0,T^{\prime}]$ and $\omega_{\psi}(v)=C(B^{\prime},\psi)v+C(B^{\prime},\psi)vh^{-1}(R/2v)$. Repeating the arguments from Lemma 3.2, we find a subsequence $\varepsilon_{m}\to 0$ such that for every $t$ the measures $\pi^{n,\varepsilon_{m}}_{t}$ converge weakly to a probability measure $\pi^{n}_{t}$. Moreover, the mapping $t\mapsto\pi^{n}_{t}$ is continuous with respect to the weak topology. The passage to the limit in the integral identity defining the solution $\pi^{n,\varepsilon_{m}}_{t}$ is based on the following two observations. First, we note that the functions $a^{ij}$, $b^{i}$, $q^{im}$ are locally bounded and continuous in $x$. Second, let us consider a function $v$ such that $v$ is integrable over $\mathbb{R}^{d}\times[0,T]$ with respect to the measure $\mu_{t}(dy)\,dt$. Let $\widetilde{v}$ be a smooth function with a compact support on $\mathbb{R}^{d}\times[0,T]$. Set
$v_{\varepsilon}(y,t)=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}v(z,s)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds$
@@
$\widetilde{v}_{\varepsilon}(y,t)=\frac{1-\varepsilon}{\varepsilon\mu_{t}^{\varepsilon}(y)}\int_{t}^{t+\varepsilon}\int_{\mathbb{R}^{d}}\widetilde{v}(z,s)\omega_{\varepsilon}(y-z)\mu_{s}(dz)\,ds.$
@@
By choosing $\widetilde{v}$, one can make the right-hand side arbitrary small. Furthermore, the functions $\widetilde{v}_{\varepsilon}$ converge uniformly to the function $\widetilde{v}$ on $\mathbb{R}^{d}\times[0,T^{\prime}]$ as $\varepsilon\to 0$. Thus to prove the passage to the limit as $\varepsilon\to 0$ we replace the functions $a^{ij}$, $b^{i}$, $q^{im}u$ and $u$ in the expressions for $a^{ij}_{\varepsilon}$, $b^{i}_{\varepsilon}$, $(Qu)_{\varepsilon}^{i}$, $u_{\varepsilon}$ by functions $\widetilde{a}^{ij}$, $\widetilde{b}^{i}$, $\widetilde{q^{im}u}$ and $\widetilde{u}$ with compact supports and then we use the uniform convergence of these new expressions $\widetilde{a}^{ij}_{\varepsilon}$, $\widetilde{b}^{i}_{\varepsilon}$, $(\widetilde{Qu})_{\varepsilon}^{i}$, $\widetilde{u}_{\varepsilon}$ as $\varepsilon\to 0$. Similarly the passage to the limit is proved in [16].
@@
Below we always assume that $\omega=\{\omega_{\psi}\}$ is constructed in the statement (iii) of Lemma 3.3. Let denote by $P_{R}^{\omega}$ the set of all measures $\Pi\in P_{R}$ such that the projection of $\Pi$ on $(x,t)$ belongs to the set $\mathcal{M}_{R,M}^{\omega}(V)$. Since $P_{R}$ and $\mathcal{M}_{R,M}^{\omega}(V)$ are compact sets in the weak topology, the set $P_{R}^{\omega}$ is compact in the weak topology.
Let us consider the mapping $(\sigma,\Pi)\mapsto\mathcal{F}_{\sigma}(\Pi)$ from $\mathcal{M}_{R,M}^{\omega}(V)\times P_{R}^{\omega}$ to $\mathbb{R}$ given by the formula
$\mathcal{F}_{\sigma}(\Pi)=\int_{U\times\mathbb{R}^{d}\times[0,T]}f(u,x,t,\sigma)\Pi(dudxdt)+\int_{\mathbb{R}^{d}}g(x,\sigma)\,\mu_{T}(dx),$
@@
holds for every $(u,x,t)\in U\times\mathbb{R}^{d}\times[0,T]$. Applying Fatou’s lemma, we derive the estimate
@@
$\int_{U\times[0,T]\times\mathbb{R}^{d}}(1-\zeta_{N}(u))|f(u,x,t,\sigma^{n})|\,\Pi^{n}(dudtdx)\leq\\ \int_{|u|>N}\Bigl(C_{h}h(|u|)+C_{f}m_{\varepsilon}+C_{f}Re^{MT}\Bigr)\Lambda(dudt)+\varepsilon TRe^{MT}.$
@@
$\int_{|u|>N}\Bigl(C_{h}h(|u|)+C_{f}m_{\varepsilon}+C_{f}Re^{MT}\Bigr)\Lambda(dudt)\leq\varepsilon.$
@@
$\int_{U\times\mathbb{R}^{d}\times[0,T]}(1-\eta_{k}(x))\zeta_{N}(u)|f(u,x,t,\sigma^{n})|\,\Pi^{n}(dudxdt)\leq\\ \leq(C_{N}\widetilde{m}_{\varepsilon}+C_{N})\Pi^{n}\bigl(\{(u,x,t)\colon|x|>k\}\bigr)+\varepsilon TRe^{MT}.$
@@
$(C_{N}\widetilde{m}_{\varepsilon}+C_{N})(\Pi^{n}+\Pi)\bigl(\{(u,x,t)\colon|x|>k\}\bigr)<\varepsilon$
@@
$\Bigl|\int_{U\times\mathbb{R}^{d}\times[0,T]}f(u,x,t,\sigma^{n})\,\Pi^{n}(dudxdt)-\\ \int_{U\times\mathbb{R}^{d}\times[0,T]}\eta_{k}(x)\zeta_{N}(u)f(u,x,t,\sigma^{n})\,\Pi^{n}(dudxdt)\Bigr|\leq 2\bigl(\varepsilon+\varepsilon TRe^{MT}\bigr).$
@@
Let $\sigma\in\mathcal{M}_{R,M}(V)$ and $u_{0}\in U$. Note that
$L_{\sigma,u_{0}}V=L_{\sigma}V+\langle Qu_{0},\nabla V\rangle\leq L_{\sigma}V+h^{*}(|Q^{\top}\nabla V|)+h(|u_{0}|).$
According to the condition (H2.1), we have the estimate
$L_{\sigma,u_{0}}V(x)\leq C_{L}V(x)+C_{L}Re^{Mt}+C_{L}\beta_{V,W}(Re^{MT})Re^{MT}+h(|u_{0}|),$
where $\beta_{V,W}$ is defined in Remark 3.2. By Remark 3.4 there exists a probability solution $\mu_{t}\,dt$ to the Cauchy problem
@@
Applying the estimate from Remark 3.4 with $C=C_{L}$, $\mathcal{V}=V$ and
$\mathcal{W}=C_{L}Re^{Mt}+C_{L}\beta_{V,W}(Re^{MT})Re^{MT}+h(|u_{0}|),$
we obtain for every $t\in[0,T]$ the estimate
$\int_{\mathbb{R}^{d}}V\,d\mu_{t}\leq e^{C_{L}T}\Bigl(\|V\|_{L^{1}(\nu)}+C_{L}T\beta_{V,W}(Re^{MT})Re^{MT}+Th(|u_{0}|)\Bigr)+\frac{1}{4}Re^{Mt}.$
Here we also use the equality $M=5C_{L}$. By Remark 3.2 we have $\lim_{R\to\infty}\beta_{V,W}(R)=0$. It follows that there exists a number $R_{1}>0$ such that for every $R>R_{1}$ the estimate
$e^{C_{L}T}\Bigl(\|V\|_{L^{1}(\nu)}+C_{L}T\beta_{V,W}(Re^{MT})Re^{MT}+Th(|u_{0}|)\Bigr)\leq\frac{3R}{4}$
holds. Let $R>R_{1}$. Then the inequality
$\int_{\mathbb{R}^{d}}V\,d\mu_{t}\leq Re^{Mt}$
is fulfilled for every $t\in[0,T]$. Set
@@
Using the inequality
$\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}W(x)\,\mu_{t}(dx)+\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}W(x)\,\sigma_{t}(dx)\leq 2\beta(Re^{MT})Re^{MT}$
and applying the conditions (H3.1) and (H3.2), we obtain the estimate
@@
Thus it suffices to minimize the functional $(u,\mu)\mapsto F_{\sigma}(u,\mu)$ only over the set of pairs $(u,\mu)$ such that $\mu_{t}\,dt$ is a solution to the Cauchy problem $\partial_{t}\mu_{t}=L^{*}_{\sigma,u(x,t)}\mu_{t}$, $\mu_{0}=\nu$, and
@@
Note that the right–hand side has the form $\alpha(Re^{MT})Re^{MT}$, where $\lim_{R\to\infty}\alpha(R)=0$. By Lemma 3.1 there exists a number $R_{2}>R_{1}$ such that for all $R>R_{2}$ and all $t\in[0,T]$ the inequalities
$\int_{\mathbb{R}^{d}}V\,d\mu_{t}\leq Re^{Mt},\quad\int_{0}^{T}\int_{\mathbb{R}^{d}}h(|u(x,t)|)\,d\mu_{t}\,dt\leq\gamma R$
are fulfilled for every Borel mapping $u\colon\mathbb{R}^{d}\times[0,T]\to U$ and for every measure $\mu\in\mathcal{M}(V)$ satisfying the following conditions: 1) the measure $\mu=\mu_{t}\,dt$ is a solution to the Cauchy problem $\partial_{t}\mu=L^{*}_{\sigma,u(x,t)}\mu$, $\mu_{0}=\nu$, 2) the estimate $F_{\sigma}(u,\mu)\leq\alpha(Re^{MT})Re^{MT}$ holds. Let $R>R_{2}$. Taking into account Remark 3.5, we conclude that it suffices to minimize the functional $(u,\mu)\mapsto F_{\sigma}(u,\mu)$ only over the set of pairs $(u,\mu)$ such that $\mu_{t}\,dt$ is a solution to the Cauchy problem $\partial_{t}\mu_{t}=L^{*}_{\sigma,u(x,t)}\mu_{t}$, $\mu_{0}=\nu$, and
@@
on $U\times\mathbb{R}^{d}\times[0,T]$. If $\mu$ is a solution to the Cauchy problem $\partial_{t}\mu_{t}=L_{\sigma,u(x,t)}^{*}\mu_{t}$, $\mu_{0}=\nu$, then $\Pi$ belongs to the set $S_{R}(\sigma)$ (see the definition before Remark 3.8).
@@
$\mathcal{F}_{\sigma}(\Pi)=\int_{U\times\mathbb{R}^{d}\times[0,T]}f(u,x,t,\sigma)\Pi(dudxdt)+\int_{\mathbb{R}^{d}}g(x,\sigma)\,\mu_{T}(dx).$
@@
the measure $\widetilde{\Pi}$ belongs to the set $S_{R}(\sigma)$. Thus we can minimize the functional $\mathcal{F}_{\sigma}(\Pi)$ on $S_{R}(\sigma)$ instead of $F_{\sigma}(u,\mu)$. Let us take a number $R>R_{2}$ such that for $S_{R}(\sigma)$ the statements of Lemma 3.3 and Lemma 3.4 hold.
Denote by $M_{R}(\sigma)$ the set of all minimizers of the functional $\mathcal{F}_{\sigma}(\Pi)$ over $S_{R}(\sigma)$. By Lemma 3.5 the mapping $\Pi\mapsto\mathcal{F}_{\sigma}(\Pi)$ is lower semi-continuous. By Lemma 3.3 the set $S_{R}(\sigma)$ is compact. Thus $M_{R}(\sigma)$ is compact and contains at least one element. Moreover, the set $M_{R}(\sigma)$ is convex since the mapping $\Pi\mapsto\mathcal{F}_{\sigma}(\Pi)$ is linear.
@@
According to Lemma 3.3, for every $\Pi\in S_{R}(\sigma)$ the projection of $\Pi$ on $(x,t)$ belongs to the set $\mathcal{M}_{R,M}^{\omega}(V)$, where $\omega=\{\omega_{\psi}\}$ does not depend on $\sigma$. To prove Theorem 2.1 it suffices to find a measure $\mu\in\mathcal{M}_{R,M}^{\omega}(V)$ such that there exists a minimizer $\Pi$ of the functional $\mathcal{F}_{\mu}(\Pi)$ on the set $S_{R}(\mu)$ and the projection of $\Pi$ on $(x,t)$ is equal to the measure $\mu$. Thus the measure $\mu$ is a fixed point of the mapping $\sigma\mapsto\Phi(\sigma)$, where $\Phi(\sigma)$ is the set of all projections of measures $\Pi\in M_{R}(\sigma)$ on $(x,t)$. Let $e_{x,t}(u,x,t)=(x,t)$. Denote by $\Pi\circ e_{x,t}^{-1}$ the projection of $\Pi$ on $(x,t)$. Note that the mapping $\Pi\mapsto\Pi\circ e_{x,t}^{-1}$ is linear and continuous with respect to the weak topology. It follows that the set $\Phi(\sigma)$ is nonempty, convex and compact in the weak topology. Recall that $\mathcal{M}_{R,M}^{\omega}(V)$ is convex (see Lemma 3.2).
The existence of a fixed point of $\Phi$ is based on the Kakutani–Ky Fan theorem (see, for instance, [19]): if a multivalued mapping $\Psi$ from a convex compact set $\mathcal{C}$ in a locally convex space to the set of non-empty convex compact subsets of $\mathcal{C}$ has a closed graph, then there exists a point $p$ such that $p\in\Psi(p)$.
Thus we need to verify that our mapping $\Phi$ has a closed graph. Assume that $\mu^{n}\in\Phi(\sigma^{n})$, $\sigma^{n}\to\sigma$ and $\mu^{n}\to\mu$. Let $\mu^{n}=\Pi^{n}\circ e_{x,t}^{-1}$, where $\Pi^{n}\in M_{R}(\sigma^{n})$. Since $\Pi^{n}\in P_{R}^{\omega}$ and $P_{R}^{\omega}$ is a compact set (see the definition before Lemma 3.5), there exists a subsequence $\Pi^{n_{k}}$ such that the measures $\Pi^{n_{k}}$ converge weakly to a measure $\Pi\in P_{R}^{\omega}$. Note that $\mu=\Pi\circ e_{x,t}^{-1}$. In what follows we assume that the original sequence $\Pi^{n}$ converges weakly to $\Pi$. We need to prove that $\Pi\in M_{R}(\sigma)$.
@@
Arguing as in the part (II), we get $\widetilde{\Gamma}\in S_{R}(\sigma)$ and $\mathcal{F}_{\sigma}(\widetilde{\Gamma})\leq\mathcal{F}_{\sigma}(\Gamma)$.
@@
Let us consider the function $u$ and the measure $\mu$ constructed in Theorem 2.1. The measure $\mu$ is a probability solution to the Cauchy problem $\partial_{t}\mu_{t}=L_{\mu,u(x,t)}^{*}\mu_{t}$, $\mu_{0}=\nu$. Moreover, we have $V\in L^{1}(\mu)$. According to the condition (H2.3), the coefficients of $L_{\mu,u(x,t)}$ are integrable over $\mathbb{R}^{d}\times[0,T]$ with respect to the measure $\mu_{t}\,dt$. Applying the superposition principle (see, for instance [16], [44]) to the measure $\mu_{t}\,dt$, we obtain a probability measure $P$ on the space $C([0,T],\mathbb{R}^{d})$ such that $\mu_{t}=P\circ e_{t}^{-1}$, where $e_{t}(\omega)=\omega(t)$, and for every function $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ the process
@@
is a martingale with respect to the measure $P$ and the natural filtration $\mathcal{F}_{t}=\sigma(\omega(s),s\leq t)$. By Proposition 2.1 from [36, Chapter 4] there exists a filtered probability space $(\Omega,\mathcal{F}_{t},\mathcal{P})$ supporting a $\mathcal{F}_{t}$–Brownian motion $W$ and a $\mathcal{F}_{t}$–adapted process $X$ such that
@@
and $\mathcal{P}\circ X_{t}^{-1}=\mu_{t}$ for all $t\in[0,T]$. Thus we take the pair $(u,\mu)$ satisfying the assertions (i), (ii) and (iii) of Theorem 2.1, apply the superposition principle and obtain the processes $X_{t}$ and $u(t,X_{t})$ satisfying the assertions (i) and (ii) of Corollary 2.1. Let us verify whether assertion (iii) of Corollary 2.1 holds for $\mu_{t}$, $X_{t}$ and $u(t,X_{t})$.